A car is moving 14.4 m/s when it hits the brakes. It slow from 78.8 m, which take 9.92s. It then accelerates at 1.83 m/s 2 for 4.50 s. What is the total displacement?
1
Expert's answer
2020-05-08T16:11:37-0400
"0.5(v_1+v_0)t_1=s_1\to 0.5(v_1+14.4)9.92=78.8"

"v_1=1.487\frac{m}{s}"

"s=s_1+v_1t_2+0.5a_2t_2^2"

"s=78.8+(1.487)(4.5)+0.5(1.83)(4.5)^2=104\ m"